# A Unifying Framework for Parallelizing Sequential Models with Linear Dynamical Systems

### v1

@@ 3.2 Case Study #2: Parallelizing RNNs with quasi-Newton iterations
Figure 2: First-order iterations can excel for complicated dynamics. We evaluate GRUs with random parameter initialization for different sequence lengths $T$ and hidden state sizes $D$. A: The nonlinear dynamics of a GRU, following Feng et al. (2024), where $x_{t}$ is the hidden state, $u_{t}$ is the input, and the notation $\mathrm{Linear}[\cdot,\cdot]$ indicates a linear readout from the concatenation of two vectors. B: A representative Jacobian matrix $\nicefrac{{\partial f_{t}}}{{\partial x}}$ from a GRU trajectory, which is not well approximated by the identity matrix. C: For each fixed-point method and a range of sequence lengths, $T$, and state sizes, $D$, we compute the median (over ten random seeds) number of fixed-point iterations to converge (top row) and the median wall-clock time (bottom row). While first-order methods yield order-of-magnitude speed-ups over sequential evaluation, Picard iterations take nearly $T$ iterations to converge.
We next consider a task where first-order fixed-point iterations like quasi-Newton and Newton iterations do well, but zeroth-order methods like Picard do not. This task is parallelizing recurrent neural networks (RNNs), like the Gated Recurrent Unit or GRU (Cho et al., 2014).
We show the results of a simple experiment in Figure˜2. We evaluate GRUs with random parameter initialization for different hidden dimension sizes $D$ and sequence lengths $T$ using sequential evaluation as well as zeroth-order (Picard) and first-order (Newton and quasi-Newton) iterations. Because the GRU in general has complicated dynamics that are not well-approximated by the identity, we observe that Picard iterations take a prohibitively large number of fixed-point iterations to converge (effectively equal to the sequence length). Therefore, Picard iterations are not suitable for these classes of dynamical systems. By contrast, first-order fixed-point iterations can yield orders of magnitude speed-ups over sequential evaluation.
@@
3.3 Case Study #3: Parallelizing discretized Langevin diffusion with Picard iterations
@@
Figure 3: Zeroth-order fixed-point iterations work well when the Jacobian is close to the identity. We evaluate Langevin dynamics for a potential $\phi$. A: The nonlinear dynamics of Langevin dynamics for a potential $\phi$ and step size $\epsilon$, where $x_{t}$ is the state and $w_{t}$ is Gaussian noise. B: The Jacobian for Langevin dynamics is well-approximated by the identity matrix, especially for small step size $\epsilon=$1\text{\times}{10}^{-5}$\$. C: We evaluate Langevin dynamics for larger dimensions, plotting the median of 10 random seeds. Picard iterations converge in a relatively small number of fixed-point iterations, as the Jacobian is well-approximated by an identity matrix. Here, zeroth- and first-order methods are often comparable in wall-clock time. The missing Newton iteration points indicate the GPU ran out of memory.
@@
Based on Table˜1, we expect that if the Jacobian of the dynamics function is well-approximated by the identity matrix, then Picard should converge relatively quickly and at considerably lower cost. A canonical example of such a system comes from Langevin dynamics (Langevin, 1908; Friedman, 2022), which are a workhorse for MCMC (Besag, 1994) and motivated the development of score-matching methods (Song & Ermon, 2019), which are closely related to diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b).
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$. Since the Jacobian is, $\tfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t})=I_{D}-\epsilon\nabla^{2}\phi(x_{t})$, it follows that the identity approximation should work well with small step-sizes, $\epsilon$. More generally, the identity approximation tends to be well-suited to problems where a differential equation is discretized with small step sizes, such as when as sampling from diffusion models (Holderrieth & Erives, 2025). In these settings, the heuristic guidance from Table˜1 suggests that Picard iterations will enjoy reasonably fast convergence.
We illustrate this heuristic guidance with a simple experiment shown in Figure˜3. We simulate Langevin dynamics on a potential $\phi$ given by the negative log probability of the mixture of two anisotropic Gaussians. In this setting, Picard iterations take far fewer than $T$ iterations to converge and can be faster than sequential evaluation. We note that quasi-Newton iterations, which include information only about the diagonal of the Jacobian of the dynamics, appear to have comparable wall-clock time, by virtue of taking fewer iterations to converge (though each fixed-point iteration involves more work).
Whether fixed-point iterations are faster than sequential evaluation also depends on memory utilization. For example, Shih et al. (2023) demonstrated wall-clock speed-ups when using Picard iterations for sampling from a diffusion model using a “sliding window” to only evaluate chunks of the sequence length where the parallel scan algorithm can fit in memory. We discuss these considerations further in Section˜C.4. However, this simple experiment simply aims to show that there are settings where zeroth-order fixed-point iterations can outperform sequential evaluation or first-order fixed-point iterations. We leave more hardware-aware implementations for future work.
@@ 4 Related Work
In this paper we unify prominent fixed-point methods for the parallel evaluation of sequences in the language of linear dynamical systems. While many papers have employed different fixed-point iterations for different problems in machine learning — Lim et al. (2024) and Danieli et al. (2023) using Newton iterations, Gonzalez et al. (2024) using quasi-Newton iterations, Shih et al. (2023) using Picard iterations, and Song et al. (2021a) using Jacobi iterations, among other works — to the best of our knowledge no one has explicitly unified these different methods in the language of linear dynamical systems.
@@
While connections between Newton’s method and Picard iterations have been made before outside of the machine learning literature, our contribution is the tight coupling of these methods to LDSs in the context of parallel evaluation of nonlinear sequences. Ortega & Rheinboldt (1970, Ch. 7) considered the problem of solving a nonlinear equation $F(\mathbf{x})=0$. They showed that Newton and Picard iterations are special cases of what they call parallel-chord methods [^3] The term “parallel” here does not have to do with parallelized computation over a sequence length, but rather comes from a geometric perspective on Newton’s method: solve a nonlinear equation by iteratively solving a linear approximation to it, i.e. forming a “chord” that is “parallel” to the nonlinear function at our current guess for its root., where each iterate is given by
$\mathbf{x}^{(i+1)}=\mathbf{x}^{(i)}-T^{-1}F(\mathbf{x}^{(i)}),$
for some matrix $T$, and they proved convergence rates for these methods. We discuss the relationship between the unifying frameworks put forward in Ortega & Rheinboldt (1970) and in our paper at greater length in Section˜D.1. The primary difference is that by focusing on the setting of nonlinear sequence evaluation, we bring into greater focus the role of the Jacobian of the dynamics function. Moreover, by unifying fixed-point iterations in the language of LDSs, we emphasize their parallelizability over the sequence length using the parallel scan (Blelloch, 1990).
@@
In this note, we focus on Jacobi, Picard, and Newton iterations because of their prominence (Song et al., 2021a; Shih et al., 2023; Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024; Grazzi & Zanella, 2025; Zoltowski et al., 2025) and their relationship to LDSs, as listed in Table˜1. However, there is a wide literature on iterative solvers (Ortega & Rheinboldt, 1970; Young, 2014). Many of these other methods can also be parallelized over the sequence length, or provide a mixture of parallel and sequential computation. For example, Song et al. (2021a) considers Gauss-Seidel iterations. Although Gauss-Seidel iterations reduce to sequential evaluation when applied to Markovian processes, Song et al. (2021a) also emphasize how the structure of the problem and hardware considerations dictate the optimal mixture of parallel and sequential computation. Parareal iterations mix parallel and sequential computation by applying parallelization at multiple length scales, and have also been used to parallelize diffusion models (Selvam et al., 2024). Tang et al. (2024) also parallelized diffusion models using both a generalization of Jacobi iterations, as well as Anderson acceleration (Anderson, 1965; Walker & Ni, 2011), which they modify to be a form of quasi-Newton. We discuss other fixed-point methods in further detail in Section˜D.2.
@@ 5 Discussion
This work unified a variety of approaches for parallelizing recursions via fixed-point iterations — including zeroth-order methods like Jacobi and Picard iterations as well as first-order methods like Newton and quasi-Newton iterations — under a common framework. In each case, the iterates reduce to evaluating an appropriately constructed linear dynamical system, which approximates the nonlinear recursion of interest. Moreover, we have demonstrated how this unifying framework provides insight into which different problems in machine learning are likely to benefit from which types of fixed-point iterations.
Clarifying the relationships and properties of these approaches through the lens of linear dynamical systems also suggests promising areas for future study. One clear direction of future work is to explore additional approaches for exploiting problem-specific structure. For example, an intermediate between Picard and quasi-Newton methods is a scaled identity approximation, $A_{t+1}=a_{t+1}I_{D}$. If we had prior knowledge on the appropriate scaling factors, $a_{t+1}\in\mathbb{R}$, we could avoid computing any Jacobian-vector product evaluations. More generally, there exist other groups of structured matrices with compact representations that are closed under composition such that a parallel evaluation of the LDS would be computationally efficient. Examples include permutation matrices, block-diagonal matrices, and block matrices where each sub-block is diagonal, among others. Future work should enumerate these use cases and investigate problem-specific applications where they are appropriate. One example application is for more efficient parallelization of the group word problem using a compact representation of permutation matrices.
In conclusion, understanding the shared backbone of these fixed-point methods can also give practitioners guidance about which methods to use for which problems. As parallel evaluation of seemingly sequential processes becomes increasingly important in machine learning, these insights may provide valuable guidance to the field.
@@ Appendix A Author Contributions and Acknowledgments
XG, EKB, and SWL conceived the project. XG and SWL proved Proposition˜1. XG and JWL proved Proposition˜3. HDL developed, ran, and wrote up the state tracking experiments. XG, EKB, HDL, JWL, and KAW all wrote substantial code in various exploratory aspects of the project. DZ and CR contributed important insights about the relationships between algorithms and computational efficiency. All authors made important contributions to conceptualization, brainstorming, exploration, writing, and editing. XG, EKB, and SWL wrote the manuscript with input from all authors. SWL supervised the project.
@@ Acknowledgments
We thank Francois Chaubard, Libby Zhang, and Henry Smith for helpful feedback. X.G. would like to acknowledge support from the Walter Byers Graduate Scholarship from the NCAA. J.W.L. was supported by the Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. S.W.L. was supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation. We thank Stanford University and the Stanford Research Computing Center for providing computational resources and support. Additional computations were performed on Marlowe (Kapfer et al., 2025), Stanford University’s GPU-based Computational Instrument, supported by Stanford Data Science and Stanford Research Computing. The authors have no competing interests to declare.

### v2

@@ 4.2 Case Study #2: Picard iterations struggle to parallelize RNNs
Figure 3: Picard iterations struggle to parallelize RNNs. We evaluate GRUs with random parameter initialization for different sequence lengths $T$ and hidden state sizes $D$. A: The nonlinear dynamics of a GRU, following Feng et al. (2024), where $x_{t}$ is the hidden state, $u_{t}$ is the input, and the notation $\mathrm{Linear}[\cdot,\cdot]$ indicates a linear readout from the concatenation of two vectors. B: A representative Jacobian matrix $\nicefrac{{\partial f_{t}}}{{\partial x}}$ from a GRU trajectory, which is not well approximated by the identity matrix. C: For each fixed-point method and a range of sequence lengths, $T$, and state sizes, $D$, we compute the median (over ten random seeds) number of fixed-point iterations to converge (top row) and the median wall-clock time (bottom row). Picard iterations take nearly $T$ iterations to converge, while the other fixed point methods yield order-of-magnitude speed-ups over sequential evaluation
Next, we consider a task where Picard iterations struggle and other fixed-point methods excel, namely, parallelizing recurrent neural networks (RNNs) like the Gated Recurrent Unit (GRU; Cho et al., 2014). We evaluate GRUs with random parameter initialization for different hidden dimension sizes $D$ and sequence lengths $T$ using sequential evaluation as well as fixed-point iterations. Figure˜3B shows that at initialization the Jacobian has small entries (on the order of $0.1$). In this regime, $\mathrm{diff}(\mathcal{A}_{J})$ and $\mathrm{diff}(\mathcal{A}_{QN})$ are less than one, while $\mathrm{diff}(\mathcal{A}_{P})$ is greater than one. (See also, Figure˜8 in Section˜E.3.) Thus, as expected, Newton, quasi-Newton, and Jacobi iterations excel in this setting, while Picard iterations converge prohibitively slowly, as shown in Figure˜3C.
@@
4.3 Case Study #3: Jacobi iterations struggle to parallelize discretized Langevin diffusion
@@
Figure 4: Jacobi iterations struggle when the dynamics Jacobian is close to the identity. We evaluate Langevin dynamics for a potential $\phi$. A: The nonlinear dynamics of Langevin dynamics for a potential $\phi$ and step size $\epsilon$, where $x_{t}$ is the state and $w_{t}$ is Gaussian noise. B: The Jacobian for Langevin dynamics is well-approximated by the identity matrix, especially for small step size $\epsilon=$1\text{\times}{10}^{-5}$\$. C: We evaluate Langevin dynamics for larger dimensions, plotting the median of 10 random seeds. Jacobi iteration consistently take $T$ steps and are always slower than sequential, while the other fixed-point methods converge in fewer $T$ steps and can be faster than sequential. The missing Newton iteration points indicate the GPU ran out of memory.
@@
Based on the theoretical analysis presented in Proposition˜3, we expect that if the Jacobian of the dynamics function is well-approximated by the identity matrix, then Picard should converge relatively quickly and at considerably lower cost, especially when compared to the other zeroth-order method of Jacobi iterations. A canonical example of such a system where the dynamics are close to identity comes from a discretization of Langevin dynamics (Langevin, 1908; Friedman, 2022). Langevin dynamics are a workhorse for MCMC (Besag, 1994) and motivated the development of score-matching methods (Song and Ermon, 2019), which are closely related to diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b).
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$. Since the Jacobian is, $\tfrac{\partial f_{t+1}}{\partial x_{t}}(x_{t})=I_{D}-\epsilon\nabla^{2}\phi(x_{t})$, it follows that the identity approximation should work well with small step-sizes $\epsilon$. More generally, the identity approximation tends to be well-suited to problems where a differential equation is discretized with small step sizes, such as when sampling from diffusion models (Holderrieth and Erives, 2025; Lai et al., 2025). In fact, simply by observing the structure of the Jacobian in Panel B of Figure˜4, we observe that $\mathrm{diff}(\cdot)$ operator for Newton, quasi-Newton, and Picard iterations in this setting will be close to zero, while $\mathrm{diff}(\mathcal{A}_{J})$ will be close to one. Therefore, based on our analysis in Proposition˜3, we hypothesize that the other fixed-point methods should dramatically outperform Jacobi iterations in this setting.
We test this hypothesis with a simple experiment shown in Figure˜4. We simulate Langevin dynamics on a potential $\phi$ given by the negative log probability of the mixture of two anisotropic Gaussians. In this setting, Picard iterations take far fewer than $T$ iterations to converge and can be faster than sequential evaluation. We note that quasi-Newton iterations, which include information only about the diagonal of the Jacobian of the dynamics, appear to have comparable wall-clock time, by virtue of taking fewer iterations to converge (though each fixed-point iteration involves more work).
Whether fixed-point iterations are faster than sequential evaluation also depends on memory utilization. For example, Shih et al. (2023) and Lu et al. (2025) demonstrated wall-clock speed-ups when using Picard iterations for sampling from a diffusion model using a “sliding window” to only evaluate chunks of the sequence length where the parallel scan algorithm can fit in memory. We discuss these considerations further in Section˜E.5.
@@ 5 Related Work
In this paper we unify prominent fixed-point methods for the parallel evaluation of sequences in the language of linear dynamical systems. While many papers have employed different fixed-point iterations for different problems in machine learning—Lim et al. (2024), Danieli et al. (2023), and Danieli et al. (2026) using Newton iterations, Tang et al. (2024) and Gonzalez et al. (2024) using quasi-Newton iterations, Shih et al. (2023) using Picard iterations, and Song et al. (2021a) using Jacobi iterations, among other works—to the best of our knowledge no one in the machine learning community has explicitly unified these different methods in the language of linear dynamical systems. It is important to note, however, that Gander and Vandewalle (2007) propose an expansive notion of parareal iterations which include these LDSs as a special case. However, Gander and Vandewalle (2007) focus on solving continuous time differential equations and do not discuss the parallel scan that has become a workhorse for the application of parallel-in-time methods on GPUs in machine learning.
@@
While connections between Newton’s method and Picard iterations have been made before outside of the machine learning literature, our contribution is the tight coupling of these methods to LDSs in the context of parallel evaluation of nonlinear sequences. Ortega and Rheinboldt (1970, Ch. 7) considered the problem of solving a nonlinear equation $\mathbf{F}(\mathbf{x})=\mathbf{0}$. They showed that Newton and Picard iterations are special cases of general iterative methods where each iterate is given by
$\mathbf{x}^{(i+1)}=\mathbf{x}^{(i)}-\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})^{-1}\mathbf{F}(\mathbf{x}^{(i)}),$ (17)
for some matrix $\widetilde{\mathbf{J}}(\mathbf{x}^{(i)})$. We discuss the relationship between the unifying frameworks put forward in Ortega and Rheinboldt (1970) and in our paper at greater length in Section˜F.1. The primary difference is that by focusing on the setting of nonlinear sequence evaluation, we bring into greater focus the role of the Jacobian of the dynamics function. Moreover, by unifying fixed-point iterations in the language of LDSs, we emphasize their parallelizability over the sequence length using the parallel scan (Blelloch, 1990).
Convergence rates of fixed-point methods
In the context of analysis of fixed-point methods in general, there is a broad literature (Ortega and Rheinboldt, 1970; Young, 2014) on the convergence rates of different fixed-point methods. For example, Ortega and Rheinboldt (1970) also proved convergence rates for iterative methods of the form in eq.˜17. Though their methods have much in common with the proof techniques used to prove Proposition˜3 of this paper, their provided results are actually trivial in the setting considered in this paper. Part of the reason [^4] We elaborate in Section F.1. for the inapplicability of the convergence results from Ortega and Rheinboldt (1970) to our paper is that Ortega and Rheinboldt (1970) consider the asymptotic setting, while it has been firmly established that in the particular setting considered in this paper, Jacobi, Picard, quasi-Newton, and Newton iterations all globally converge in at most $T$ iterations (Bellen and Zennaro, 1989; Shih et al., 2023; Tang et al., 2024; Gonzalez et al., 2024). For moving beyond this worst-case analysis, Gonzalez et al. (2025) uses an optimization perspective to show that the difficulty of parallelizing a dynamical system is directly related to the stability of the system, which can be thought of as the “average” spectral norm of $\nicefrac{{\partial f_{t+1}}}{{\partial x_{t}}}$. Proposition 4 of Lu et al. (2025) nicely presents the foundations of the convergence analysis we present in Proposition˜3 of this paper. However, we extend their work by applying it to a wider variety of fixed-point methods, explicitly bounding many quantities of interest, and demonstrating its relevance in simulation.
@@
In this note, we focus on Jacobi, Picard, and Newton iterations because of their prominence (Song et al., 2019; 2021a; Shih et al., 2023; Danieli et al., 2023; Lim et al., 2024; Gonzalez et al., 2024; Grazzi and Zanella, 2025; Zoltowski et al., 2025; Farsang and Grosu, 2025; Danieli et al., 2026; Iacob et al., 2025) and their relationship to LDSs, as listed in Table˜1. However, there is a wide literature on iterative solvers (Ortega and Rheinboldt, 1970; Young, 2014). Many of these other methods can also be parallelized over the sequence length, or provide a mixture of parallel and sequential computation. For example, Horton et al. (1995) apply the parallel associative scan (which they call parallel cyclic reduction) for parallel-in-time differential equation solvers, though they focus on Jacobi and Gauss-Seidel iterations. In discrete time, Naumov (2017) shows how evaluating Markov chains can be cast as a system of nonlinear equations and discusses many techniques from numerical analysis for solving them, again focusing on Jacobi and Gauss-Seidel; Song et al. (2021a) extend this program with many deep learning experiments. Although Gauss-Seidel iterations reduce to sequential evaluation when applied to Markovian processes, Song et al. (2021a) also emphasize how the structure of the problem and hardware considerations dictate the optimal mixture of parallel and sequential computation. Parareal iterations (Lions et al., 2001) mix parallel and sequential computation by applying parallelization at multiple length scales, and have also been used to parallelize diffusion models (Selvam et al., 2024). Tang et al. (2024) also parallelize diffusion models using both a generalization of Jacobi iterations, as well as Anderson acceleration (Anderson, 1965; Walker and Ni, 2011), which they modify to be a form of quasi-Newton. We discuss other fixed-point methods based on trust-region techniques in Section˜F.2.
@@ 6 Discussion
This work unified a variety of approaches for parallelizing recursions via fixed-point iterations—including zeroth-order methods like Jacobi and Picard iterations as well as first-order methods like Newton and quasi-Newton iterations—under a common framework. In each case, the iterates reduce to evaluating an appropriately constructed linear dynamical system, which approximates the nonlinear recursion of interest. This unifying framework provides insight into which different problems in machine learning are likely to benefit from which types of fixed-point iterations. In particular, we demonstrate that the structure of the Jacobian matrix of the dynamics function plays a key role in determining which fixed-point method to use.
For this reason, understanding the structure of the Jacobian of the dynamics function is important for using our framework. Fortunately, there are many problems where the structure of the Jacobian matrix is known in advance. As we showed in Section˜4.1, the group word problem can always be simulated with permutation matrices for its dynamics. As we showed in Section˜4.3, discretized roll-outs from differential equations, used in sampling from diffusion models and rolling out neural ODEs, have $\nicefrac{{\partial f}}{{\partial x}}$ equal to the identity matrix plus a correction term proportional to the discretization step-size. Moreover, as shown in Zoltowski et al. (2025), the dynamics of position and momenta variables in Hamiltonian Monte Carlo (HMC) results in banded matrices. Furthermore, in sequence modeling, one can design a recurrent neural network to have Jacobians with desired structure (Farsang and Grosu, 2025; Zattra et al., 2025; Danieli et al., 2026). Finally, if there is truly no analytic information about the Jacobian in advance, its structure can be probed with finite-difference methods.
Future directions
Clarifying the relationships and properties of these approaches through the lens of linear dynamical systems also suggests promising areas for future study. One clear direction of future work is to explore additional approaches for exploiting problem-specific structure, using our unifying framework to develop new fixed-point iterations. For example, an intermediate between Picard and quasi-Newton methods is a scaled identity approximation, $\tilde{A}_{t+1}=a_{t+1}I_{D}$. If we had prior knowledge on the appropriate scaling factors, $a_{t+1}\in\mathbb{R}$, we could avoid computing any Jacobian-vector product evaluations. More generally, there exist other groups of structured matrices with compact representations that are closed under composition such that a parallel evaluation of the LDS would be computationally efficient. Examples include permutation matrices, block-diagonal matrices, and block matrices where each sub-block is diagonal, among others. Future work should enumerate these use cases and investigate problem-specific applications where they are appropriate. One example application is for more efficient parallelization of the group word problem using a compact representation of permutation matrices, as was done in concurrent work by Terzić et al. (2025).
Understanding the shared backbone of these fixed-point methods can give practitioners guidance about which methods to use for which problems. As parallel evaluation of seemingly sequential processes becomes increasingly important in machine learning, these insights may provide valuable guidance to the field.
@@ Appendix A Author Contributions and Acknowledgments
XG, EKB, and SWL conceived the project. XG and SWL proved Proposition˜1 and Proposition˜2. XG provided most of the content of Section˜3, with input from all authors. XG and JWL proved Proposition˜4. HDL developed, ran, and wrote up the state tracking experiments. XG, EKB, HDL, JWL, and KAW all wrote substantial code in various exploratory aspects of the project. DZ and CR contributed important insights about the relationships between algorithms and computational efficiency. LK contributed important insights about rates of convergence and the importance of stability. All authors made important contributions to conceptualization, brainstorming, exploration, writing, and editing. XG, EKB, and SWL wrote the manuscript with input from all authors. SWL supervised the project.
@@ Acknowledgments
We thank Nicolas Zucchet, Francois Chaubard, Libby Zhang, Henry Smith, and Ian Christopher Tanoh, as well as the anonymous reviewers on Open Review, for helpful feedback. X.G. acknowledges support from the Walter Byers Graduate Scholarship from the NCAA. J.W.L. was supported by the Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. S.W.L. was supported by fellowships from the Simons Collaboration on the Global Brain, the Alfred P. Sloan Foundation, and the McKnight Foundation. We thank Stanford University and the Stanford Research Computing Center for providing computational resources and support. Additional computations were performed on Marlowe (Kapfer et al., 2025), Stanford University’s GPU-based Computational Instrument, supported by Stanford Data Science and Stanford Research Computing. The authors have no competing interests.